# ArchesClimate: Probabilistic Decadal Ensemble Generation With Flow Matching

### v1

@@ 2.4 Inference
Once both the deterministic and generative models are trained, ArchesClimate auto-regressively generates sequential states. To generate a state, the model needs to move from Gaussian noise $\epsilon$ to the data distribution of the following step. ArchesClimate takes $M\in\mathbb{N}$ FM steps during inference to go from $r_{t+\delta,0}$ to $r_{t+\delta,S}$ where $M$ is a hyperparameter set at inference time. At each FM inference step $m$, the model outputs $r_{t+\delta,\psi_{m+1}}$ with ArchesClimate until it reaches $r_{t+\delta,S}$ where $\psi$ consists of evenly spaced values from 0 to S with a step size of $\frac{S}{M}$.
$r_{t+\delta,\psi_{m+1}}=r_{t_{+}\delta,\psi_{m}}+(\psi_{m+1}-\psi_{m})g_{\theta}(r_{t_{+}\delta,\psi_{m}}|f_{\theta}(X_{t}),X_{t-\delta})$ (4)
@@
Each step in the sampling process can be interpreted as taking a small step $(\psi_{m+1}-\psi_{m})$ in the direction given by the vector field generated with $g_{\theta}$, which is then added to the current state $r_{t_{+}\delta,\psi_{m}}$.
@@
Once ArchesClimate has followed the probability path to the data distribution, it can take $r_{t+\delta}$ and use Equation 3 to build the full next state. This is then be used as input to generate subsequent states. See Figure 3 for a visualization of this process. To initialize ArchesClimate for inference, the model begins with states taken from IPSL-DCPP. To generate states for the years 1969-1979, we will initialize ArchesClimate with January 1969 and February 1969 from a random ensemble member of IPSL-DCPP.
@@
Figure 2: Deterministic and Generative training schemes for ArchesClimate. It is necessary to have full trained $f_{\theta}$ before training $g_{\theta}$.
@@
Figure 3: Sampling with ArchesClimate. Initial states and noise are given to $g_{\theta}$ and slowly shift from noise to the data distribution. The combined result of $f_{\theta}$ and $g_{\theta}$ are then used as input for the following timestep $t$.
@@ 3 Experiments
3.1 Experimental Setup
To measure how well ArchesClimate captures the dynamics of IPSL-DCPP, we consider the following baseline. For each target ensemble, we select 5 members of the 10 members to serve as a baseline. We then compare both the remaining 5 ensemble members of IPSL-DCPP and a 5-member ArchesClimate ensemble to the baseline.
@@ 3.1 Experimental Setup
For analyses requiring the anomaly of the predicted state, we calculate the anomaly by removing the monthly mean taken from the training period.
@@ 3.2 Evaluation Metrics and Diagnostic Tools
Continuous Ranked Probability Score (CRPS) is a metric used to evaluate the accuracy of probabilistic forecasts by measuring the difference between the predicted cumulative distribution function and the observed outcome. We use the same implementation of CRPS as in \@BBOPcite\@BAP\@BBNRasp et al. (2024)\@BBCP.
Rank Histograms are a diagnostic tool used to evaluate the consistency or calibration of an ensemble forecast by comparing it to another ensemble \@BBOPcitep\@BAP\@BBN(Hamill, 2001)\@BBCP. Using rank histograms, we evaluate whether an ensemble member generated with ArchesClimate can be interchanged with any member of the IPSL-DCPP ensemble. To calculate the rank of an ensemble member, we take the pixel-by-pixel values of the ensemble member and compare them to a target ensemble. We calculate the rank of the pixel value of the ensemble member compared to the rest of the target ensemble members at that pixel. We then take an average of the rank of each pixel across space and time. The histograms show the normalized frequency of the emulated ensemble member at that rank. We normalize the frequency by taking the mean and standard deviation of all frequencies across space and time. If the generated ensemble member can be interchanged with any member of an IPSL-DCPP ensemble, the histogram will be flat, showing an even distribution across all ranks.
@@
$\mathrm{lat}=1,\dots,N_{y},\quad\mathrm{lon}=1,\dots,N_{x},\quad\mathcal{K}=\left\{\,k\;\middle|\;1\leq k\leq\left\lfloor\frac{T}{2}\right\rfloor=60\,\right\}.$
$X(k,\mathrm{lat},\mathrm{lon})=\sum_{n=0}^{119}x(n,\mathrm{lat},\mathrm{lon})\,e^{-i2\pi kn/120}.$ (5)
@@
Pearson Correlation Coefficient (PCC) measures the strength and direction of the linear relationship between two variables. It ranges from -1 to 1, where values close to 1 indicate a strong positive linear relationship, values close to -1 indicate a strong negative linear relationship, and values near 0 suggest little to no linear correlation between the variables.
Figure 4: Ensemble means of the full state (top) and Ensemble means for anomalies (bottom) of the Tropics (20° S – 20° N, 0° – 360° E) for the years 1969-1979. The dotted lines are the maximum and minimums for each ensemble mean, with the shaded area being +/- 1 standard deviation.
@@ 3.3 Regional Accuracy of ArchesClimate in the Tropics
By training on the full state of a given timestep, ArchesClimate can learn both the signal for the seasonal cycle and the anomaly, both of which are important aspects of decadal climate analysis \@BBOPcitep\@BAP\@BBN(Smith et al., 2019)\@BBCP. The seasonal cycle is the predictable annual pattern of changes caused by Earth’s position around the Sun, while anomalies are deviations from this expected seasonal average. Looking at anomalies is important in climate science because they highlight changes relative to a long-term average, making it easier to detect trends and patterns beyond natural variability. In this experiment, we qualitatively compare the performance of ArchesClimate to IPSL-DCPP in the Tropics (20°S–20°N, 0°–360°E) for both the full state and the anomaly of the full state. We use ArchesClimate to generate a 10-member ensemble and compare it to a 10-member ensemble of IPSL-DCPP for the test period initialized at 1969 for the Tropics. Each ensemble is 120 months (10 years), and we assess the stability and the regional accuracy of the generated states from ArchesClimate. We target the Tropics as they exhibit many important dynamics at climatic timescales \@BBOPcitep\@BAP\@BBN(Wang, 2019)\@BBCP. To show the performance in both ocean and atmosphere, we select a subset of variables emulated in ArchesClimate. Sea surface temperature (tos) and sea level pressure (psl) exemplify ocean dynamics, while total positive downward flux (net_flux) shows interactions between atmosphere and ocean. Air temperature at 700 hPa (ta) provides an example of atmospheric dynamics.
@@
Figure 5: Variance (top) and CRPS (bottom) for different training schemes for the decade 1969-1979. residual-flow is described in Section 2.3. full-flow trains without any deterministic model. full-deterministic uses only the deterministic model to make predictions. The dotted line represents the 5-member IPSL-DCPP baseline.
@@ 3.4 Comparison of Training Schemes
ArchesClimate uses a combination of deterministic and generative models. To assess the training scheme outlined in Section 2.3 (hereby referred to as residual-flow), we compare it to several other schemes. First, we compare to a model where the residuals (as defined in Section 2.3) are trained with Denoising Diffusion Probabilistic Models (DDPM) instead of FM and call this residual-ddpm. DDPM gradually adds Gaussian noise to data, then learns to reverse this corruption to generate samples \@BBOPcitep\@BAP\@BBN(Ho et al., 2020)\@BBCP. We then make predictions using only a deterministic model and only a probabilistic model, named full-deterministic and full-flow, respectively. We compare these methods to our baseline outlined in Section 3.1.
As residual-ddpm is unable to generate stable states for longer than approximately 20 months, we do not include it in Figure 5. We see in Figure 5 that full-deterministic exhibits low variance compared to the other methods. Full-flow has similar variance to IPSL-DCPP but is less accurate than residual-flow across all variables. By combining the probabilistic and deterministic components (as shown in residual-flow), ArchesClimate makes much more accurate predictions and generates more variance than full-deterministic. We can see in tos that both full-deterministic and full-flow have high CRPS scores compared to IPSL-DCPP, but residual-flow has a similar CRPS score to IPSL-DCPP. The probabilistic model, when trained in tandem with the deterministic model, provides error correction and improves accuracy on top of increasing variance. We investigate methods to increase the variance of residual-flow in Section 3.10. Because FM learns a deterministic continuous-time flow, it requires fewer integration steps than DDPM at inference, making it more efficient and less prone to error accumulation. The implementation of FM is significantly simpler than DDPM, with fewer hyperparameters and implementation details and therefore helps FM perform better than DDPM.
3.5 CRPS and Variance of ArchesClimate
By computing the average CRPS and Variance over each period, we assess how the model responds to different forcings and different initializations. In this experiment, we quantify the performance of ArchesClimate using CRPS and Variance and compare ArchesClimate to our baseline averaged over three periods: 1969-1979, 1979-1989 and 2010-2020. We add vertically-averaged potential temperature at 0-2000m (thetaot2000) and several atmospheric variables (Omega (wap), geopotential (zg) and relative humidity (hur)) to the subset of variables already used in Section 3.3 to display a wider range of dynamics in ArchesClimate.
In Table 2, the CRPS for all variables except for thetaot2000 is lower in ArchesClimate than in IPSL-DCPP. The variance is consistently higher in IPSL-DCPP across all periods. As we predict the mean state deterministically and the residuals probabilistically, it is possible to increase the variance of the initial noise during the generation of the residual to increase the variance. We explore several ways to improve variance in Section 3.10. A lower CRPS score means that ArchesClimate is better at capturing the distribution of our baseline, the remaining 5 members of the IPSL-DCPP, than the first 5 members of the IPSL-DCPP. It is not clear that a lower CRPS score is better, as we want to replicate the dynamics of the IPSL-DCPP and therefore have similar CRPS to IPSL-DCPP. This ambiguity in CRPS necessitates more qualitative analysis of the generated states from ArchesClimate. Qualitative analysis is carried out in the subsequent experiments.
Variable CRPS Variance
ArchesClimate IPSL-DCPP ArchesClimate IPSL-DCPP
thetaot2000 (°C) 0.06 0.05 0.02 0.05
tos (°C) 0.27 0.27 0.26 0.51
psl (Pa) 87.98 106.39 30529.94 109017.79
net_flux (W/m-2) 7.25 7.97 270.42 544.49
zg 700 hPa (m) 8.38 10.03 242.58 938.19
wap 700 hPa (Pa/s-1) 1.07e-02 1.14e-02 6.53e-04 9.67e-04
ta 700 hPa (K) 0.41 0.46 0.68 1.58
hur 700 hPa (%) 2.78 3.02 32.40 55.38

### v2

@@ 2.5 Inference
Once both the deterministic and generative models are trained, ArchesClimate auto-regressively generates sequential states. To generate a state, the model needs to move from Gaussian noise $\epsilon$ to the data distribution of the following step. ArchesClimate takes $M\in\mathbb{N}$ FM steps during inference to go from $r_{t+\delta,0}$ to $r_{t+\delta,S}$ where $M$ is a hyperparameter set at inference time. $M$ is a small fraction of $S$. At each FM inference step $s$, the model outputs $r_{t+\delta,\Delta_{s+1}}$ with ArchesClimate until it reaches $r_{t+\delta,S}$ where $\Delta$ consists of evenly spaced values from 0 to S with a step size of $\frac{S}{M}$:
$r_{t+\delta,\Delta_{s+1}}=r_{t_{+}\delta,\Delta_{s}}+(\Delta_{s+1}-\Delta_{s})g_{\theta}(r_{t_{+}\delta,\Delta_{s}}|f_{\theta}(X_{t+\delta}|X_{t},X_{t-\delta})$ (3)
@@
Each step in the sampling process can be interpreted as taking a small step $(\Delta_{s+1}-\Delta_{s})$ in the direction given by the vector field generated with $g_{\theta}$, which is then added to the current state $r_{t_{+}\delta,\Delta_{s}}$. To recreate $X_{t+\delta}$, we combine the output of the deterministic and generative model:
$X_{t+\delta}=f_{\theta}(X_{t+\delta}|X_{t},X_{t-\delta})+r_{t+\delta,S}$ (4)
@@
Once ArchesClimate has followed the probability path to the data distribution, it can take $r_{t+\delta,S}$, the output from $g_{\theta}$, and use Equation 4 to build the full next state. This is then passed to $f_{\theta}$ (and subsequently $g_{\theta}$) as input to generate subsequent states. See Figure 3 for a visualization of this process. To initialize ArchesClimate for inference, the model begins with states taken from IPSL-DCPP. To generate states for the years 1969-1979, we will initialize ArchesClimate with January 1969 and February 1969 from a random ensemble member of IPSL-DCPP.
@@
Figure 2: Deterministic and Generative training schemes for ArchesClimate. It is necessary to have fully trained $f_{\theta}$ before training $g_{\theta}$. $f_{\theta}$ learns a strong prior of the mean climate, while $g_{\theta}$ learns the residuals of the learned mean climate.
@@
Figure 3: Sampling with ArchesClimate. Initial states and noise are given to $g_{\theta}$ and slowly shift from noise to the data distribution over 24 inference timesteps. The combined result of $f_{\theta}$ and $g_{\theta}$ are then used as input for the following timestep $t$.
2.6 Training Details
Table 2: Hyperparameters used in ArchesClimate.
Hyperparameter Value
Learning Rate $4e^{-4}$
Inference Flow Matching Steps (M) 12
Training Flow Matching Steps (S) 1000
Optimizer AdamW
Schedule Cosine with Warmup
Warmup Steps 5000
Betas ($\beta_{1},\beta_{2}$) 0.9, 0.98
Weight Decay 0.03
Quantifying the precise computational speedup of ArchesClimate relative to a traditional Earth System Model (ESM) like IPSL-CM6A-LR remains challenging due to fundamental differences in architecture and resolution. However, a rough comparison can be made by estimating the effective ”parameter space” and normalizing for hardware differences.
The IPSL-CM6A-LR model operates on approximately $1.06\times 10^{7}$ spatial grid points (Acosta et al., 2024). While the model tracks roughly 750 prognostic variables, the computational load is unevenly distributed. According to Acosta et al. (2024), the ocean component comprises the majority of the grid ($9\times 10^{6}$ points), compared to the atmospheric component ($1\times 10^{6}$ points). Assuming the computational load is dominated by standard ocean variables (approx. 41 variables based on CMIP6 ”Omon” datasets), we estimate the effective parameter space of the IPSL model to be roughly $4.0\times 10^{8}$. In terms of throughput, IPSL-CM6A-LR completes 12 simulated years per day, consuming approximately 1,900 core-hours per simulated year (Acosta et al., 2024).
In contrast, ArchesClimate operates on a significantly reduced grid of $144\times 144\times 45$ (approx. $10^{6}$ points). This suggests ArchesClimate is roughly 1/400th the size of the IPSL model in terms of spatial complexity. Furthermore, we assume here a monthly temporal resolution for ArchesClimate and IPSL, whereas in practice the IPSL model computes daily and hourly variables, further widening the computational gap.
To compare costs, we must normalize the hardware usage. Using the approximation that one NVIDIA V100 GPU is roughly equivalent to 200 CPU cores (Kim et al., 2021), we can derive a ”core-hour” equivalent for the machine learning model. ArchesClimate simulates 10 years in 12 minutes, or 50 simulated years per wall-clock hour. Using the conversion factor of 200 cores, 1 hour of GPU time equals 200 core-hours. Therefore, $200/50$ results in a cost of 4 core-hours per simulated year for ArchesClimate.
While ArchesClimate is drastically cheaper in absolute terms (4 vs. 1,900 core-hours), this efficiency is largely a function of its reduced resolution. If we were to scale ArchesClimate linearly to match the IPSL’s parameter space (multiplying the cost by the size factor of  400), the theoretical cost would rise to 1,600 core-hours per simulated year. This implies that at equivalent resolutions, the ML approach might be only marginally less expensive than the physics-based GCM. Moreover, if higher temporal resolutions were required, the ML cost could exceed that of the GCM. However, linear scaling may be an overly pessimistic assumption, as ML techniques often leverage latent spaces to improve efficiency beyond brute-force scaling. The primary value of ArchesClimate is not necessarily raw per-parameter efficiency, but its ability to approximate IPSL dynamics—effectively distilling a complex model—without running the full simulation. The IPSL is not capable of doing reduced-complexity runs, and our technique to distill the model’s dynamics enables this reduced-complexity. Finally, total cost comparisons must account for training overhead, a factor often omitted in GCM estimates. The deterministic ArchesClimate model required 20 hours of training on 8 V100s, while the generative model required 20 hours on 4 A100s. Both physics-based and ML models require extensive tuning, adding further complexity to a direct ”price tag” comparison.
@@ 3 Experiments
3.1 Baselines
To measure how well ArchesClimate captures the dynamics of IPSL-DCPP, we consider the following baselines.
Held-out IPSL-DCPP Ensemble For each target ensemble, we select 5 members of the 10 members to serve as a baseline. We then compare both the remaining 5 ensemble members of IPSL-DCPP and a 5-member ArchesClimate ensemble to the baseline.
Pattern Scaling Baseline Following ClimateBench (Watson-Parris et al., 2022), we employ the Pattern Scaling method. Pattern scaling is a computationally efficient technique that estimates detailed regional climate changes for various future scenarios by simply multiplying a fixed spatial map of change (derived from complex models) by the projected global average temperature. Pattern scaling is known to be an effective means to predict spatial temperatures, but struggles with learning non-linear relationships such as precipitation (Tebaldi and Arblaster, 2014). We train a linear regression model using global-mean temperature to predict these spatially explicit means. Consistent with the ClimateBench framework, we provide the global mean temperature as input (Watson-Parris et al., 2022). We provide the pattern scaling method with the monthly global mean. We train the linear regression model with the same data used in ArchesClimate. We generate an ensemble by using the global mean temperature from each ensemble member in IPSL-DCPP. We use the resulting generated states to compare the CRPS and Variance with ArchesClimate and IPSL-DCPP.
@@ 3.2 Evaluation Metrics and Diagnostic Tools
Continuous Ranked Probability Score (CRPS) is a metric used to evaluate the accuracy of probabilistic forecasts by measuring the difference between the predicted cumulative distribution function and the observed outcome. We use the same implementation of CRPS as in Rasp et al. (2024) and is latitude-weighted.
Rank Histograms are a diagnostic tool used to evaluate the consistency or calibration of an ensemble forecast by comparing it to another ensemble (Hamill, 2001). Using rank histograms, we evaluate whether an ensemble member generated with ArchesClimate can be interchanged with any member of the IPSL-DCPP ensemble. To calculate the rank of an ensemble member, we take the pixel-by-pixel values of the ensemble member and compare them to a target ensemble. We calculate the rank of the pixel value of the ensemble member compared to the rest of the target ensemble members at that pixel. We then take an average of the rank of each pixel across space and time. The histograms show the normalized frequency of the emulated ensemble member at that rank. We normalize the frequency by taking the mean and standard deviation of all frequencies across space and time. If the generated ensemble member can be interchanged with any member of an IPSL-DCPP ensemble, the histogram will be flat, showing an even distribution across all ranks.
@@
$\mathrm{lat}=1,\dots,N_{y},\quad\mathrm{lon}=1,\dots,N_{x},\quad\mathcal{K}=\left\{\,k\;\middle|\;1\leq k\leq\left\lfloor\frac{T}{2}\right\rfloor=60\,\right\}$
$X(k,\mathrm{lat},\mathrm{lon})=\sum_{n=0}^{119}x(n,\mathrm{lat},\mathrm{lon})\,e^{-i2\pi kn/120}$ (5)
@@
Figure 4: Ensemble means of the full state (top) and ensemble means for anomalies (bottom) of the Tropics (20° S – 20° N, 0° – 360° E) for the test period 1969-1979. The dotted lines are the maximum and minimums for each 5-member ensemble mean, with the shaded area being +/- 1 standard deviation. Closer to the dotted line is better.
@@ 3.3 Regional Accuracy of ArchesClimate in the Tropics
By training on the full state of a given timestep, ArchesClimate can learn both the signal for the seasonal cycle and the anomaly, both of which are important aspects of decadal climate analysis (Smith et al., 2019). The seasonal cycle is the predictable annual pattern of changes caused by Earth’s position around the Sun, while anomalies are deviations from this expected seasonal average. Looking at anomalies is important in climate science because they highlight changes relative to a long-term average, making it easier to detect trends and patterns beyond natural variability. In this experiment, we qualitatively compare the performance of ArchesClimate to IPSL-DCPP in the Tropics (20°S–20°N, 0°–360°E) for both the full state and the anomaly of the full state. We use ArchesClimate to generate a 10-member ensemble and compare it to a 10-member ensemble of IPSL-DCPP for the test period initialized at 1969 for the Tropics. Each ensemble is 120 months (10 years), and we assess the stability and the regional accuracy of the generated states from ArchesClimate. We target the Tropics as they exhibit many important dynamics at climatic timescales (Wang, 2019). To show the performance in both ocean and atmosphere, we select a subset of variables emulated in ArchesClimate. Sea surface temperature (tos) and sea level pressure (psl) exemplify ocean dynamics, while total positive downward flux (net_flux) shows interactions between atmosphere and ocean. Air temperature (ta) at 700 hPa provides an example of atmospheric dynamics.
@@
Figure 5: Variance (top) and CRPS (bottom) for different training schemes for the decade 1969-1979. residual-flow is described in subsection 2.4. full-flow trains without any deterministic model. full-deterministic uses only the deterministic model to make predictions. The dotted line represents the 5-member IPSL-DCPP baseline.
@@ 3.4 Comparison of Training Schemes
ArchesClimate uses a combination of deterministic and generative models. To assess the training scheme outlined in subsection 2.4 (hereby referred to as residual-flow), we compare it to several other schemes. We make predictions using only a deterministic model and only a probabilistic model, named full-deterministic and full-flow, respectively. We compare these methods to our pattern scaling baseline outlined in subsection 3.1.
We see in Figure 5 that full-deterministic exhibits low variance compared to the other methods. Full-flow has similar variance to IPSL-DCPP but is less accurate than residual-flow across all variables. By combining the probabilistic and deterministic components (as shown in residual-flow), ArchesClimate makes much more accurate predictions and generates more variance than full-deterministic. We can see in tos that both full-deterministic and full-flow have high CRPS scores compared to IPSL-DCPP, but residual-flow has a similar CRPS score to IPSL-DCPP. The probabilistic model, when trained in tandem with the deterministic model, provides error correction and improves accuracy on top of increasing variance. We investigate methods to increase the variance of residual-flow in subsection 3.10. We outperform the pattern scaling baseline across all metrics, with the exception of temperature variance. While the numerical margin is small, this performance is notable given the methodological differences: the baseline is conditioned on perfect global mean temperatures for every month, whereas ArchesClimate performs a fully auto-regressive rollout over the 10-year period. By using pattern matching for temperature, variance in ArchesClimate could potentially be improved. This exploration is left for further work.
3.5 CRPS and Standard Deviation of ArchesClimate
We assess how the model responds to varying forcings and initializations by computing the average Continuous Ranked Probability Score (CRPS) and standard deviation over three distinct periods: 1969–1979, 1979–1989, and 2010–2020. To evaluate a wider range of dynamics, we expanded the variable subset used in subsection 3.3 to include vertically-averaged potential temperature at 0–2000m (thetaot2000), vertical velocity (wap), geopotential (zg), and relative humidity (hur). The pattern scaling baseline is derived from the monthly global mean temperature of each ensemble member, allowing us to generate a full ensemble for comparison. Further details on these metrics are provided in subsection 3.2.
As shown in Table 3, ArchesClimate achieves a lower CRPS than IPSL-DCPP for all variables except thetaot2000. However, the standard deviation is consistently higher in the IPSL-DCPP reference across all periods. Because ArchesClimate predicts the mean state deterministically and residuals probabilistically, the variance can be adjusted by increasing the initial noise amplitude during residual generation—a topic explored in subsection 3.10.
The interpretation of these scores requires nuance. A CRPS lower than the IPSL score indicates that ArchesClimate captures the distribution of the target IPSL-DCPP members more effectively than the internal IPSL-DCPP ensemble members do among themselves. Since our goal is to faithfully emulate the original model’s dynamics, a lower CRPS is not inherently better; rather, we aim for a score similar to that of the IPSL-DCPP baseline. This metric ambiguity necessitates the qualitative analysis of generated states conducted in subsequent experiments.
Finally, while the pattern scaling baseline yields a similar CRPS, it exhibits significantly lower variance than both IPSL-DCPP and ArchesClimate. This result is expected, as pattern scaling predicts values near the ensemble mean. Furthermore, unlike ArchesClimate, pattern scaling is not auto-regressive and is provided with the ground-truth global monthly means as input.
Variable CRPS Std. Dev.
ArchesClimate IPSL-DCPP Pattern Scaling ArchesClimate IPSL-DCPP Pattern Scaling
thetaot2000 (°C) 0.06 0.05 0.04 0.14 0.22 0.03
tos (°C) 0.27 0.27 0.18 0.51 0.71 0.01
psl (Pa) 87.98 106.39 95.18 174.73 330.18 8.3
net_flux (W/m-2) 7.25 7.97 7.27 16.44 23.33 0.6
zg 700 hPa (m) 8.38 10.03 9.32 15.57 30.63 2.7
wap 700 hPa (Pa/s-1) 1.07e-02 1.14e-02 1.00e-3 2.56e-02 3.11e-02 6.71e-04
ta 700 hPa (K) 0.41 0.46 0.44 0.82 1.26 0.25
hur 700 hPa (%) 2.78 3.02 2.57 5.69 7.44 0.27